# DSDNet: Raw Domain Demoiréing via Dual Color-Space Synergy

### v1

@@ Abstract.
With the rapid advancement of mobile imaging, capturing screens using smartphones has become a prevalent practice in distance learning and conference recording. However, moiré artifacts, caused by frequency aliasing between display screens and camera sensors, are further amplified by the image signal processing pipeline, leading to severe visual degradation. Existing sRGB domain demoiréing methods struggle with irreversible information loss, while recent two-stage raw domain approaches suffer from information bottlenecks and inference inefficiency. To address these limitations, we propose a single-stage raw domain demoiréing framework, Dual-Stream Demoiréing Network (DSDNet), which leverages the synergy of raw and YCbCr images to remove moiré while preserving luminance and color fidelity. Specifically, to guide luminance correction and moiré removal, we design a raw-to-YCbCr mapping pipeline and introduce the Synergic Attention with Dynamic Modulation (SADM) module. This module enriches the raw-to-sRGB conversion with cross-domain contextual features. Furthermore, to better guide color fidelity, we develop a Luminance-Chrominance Adaptive Transformer (LCAT), which decouples luminance and chrominance representations. Extensive experiments demonstrate that DSDNet outperforms state-of-the-art methods in both visual quality and quantitative evaluation, and achieves an inference speed 2.4x faster than the second-best method, highlighting its practical advantages. We provide an anonymous online demo at DSDNet.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: xxxxxccs: Computing methodologies Computational photography
@@ 1. Introduction
Recent advancements in sRGB-domain demoiréing have been driven by deep learning methods and the availability of large, paired datasets (Cheng et al., 2019; He et al., 2020). These methods have introduced innovative strategies such as frequency-domain modeling (He et al., 2019), multiscale feature fusion (Sun et al., 2018), and moiré pattern classification (Liu et al., 2018), leading to significant improvements in moiré removal. However, these sRGB-based methods (Zheng et al., 2020; Liu et al., 2020b) suffer from two critical limitations. First, they often overlook the underlying physical processes of moiré pattern formation, especially the fact that demosaicing in ISP amplifies moiré artifacts (Yue et al., 2022), as shown in Fig. 2(a). Second, by passing through the ISP pipeline, images suffer from the irreversible information loss (Jin et al., 2023), making restoring clean images more challenging.
In response to these limitations, recent research has shifted towards raw domain processing, which preserves the linear characteristics of the sensor signal and better aligns moiré pattern formation with the ISP pipeline. Although raw domain demoiréing frameworks have shown significant improvements over traditional sRGB methods, they still face inherent bottlenecks. Typically, these frameworks (Yue et al., 2022; Xu et al., 2024c) consist of two stages: the first stage learns the mapping from moiré to clean images, while the second stage learns the mapping from raw to sRGB, as shown in Fig. 2(b) and (c). These methods use cascaded subnetworks, where each stage enhances the output image from the previous one. With image-level data flow, only images are propagated forward, and each stage relies solely on the result of the prior stage. This design inherently increases the risk of information loss and error accumulation. Additionally, every subnetwork in each stage may incur information loss due to the downsampling operation. As a result, suboptimal performance arises from the lossy image-level data flow, where errors propagate, accumulate, and amplify across stages, ultimately failing to reconstruct texture details and color fidelity. Moreover, these two-stage pipelines (Song et al., 2023; Yue et al., 2023a) face challenges in real-time inference and introduce significant computational overhead. For instance, Xu et al. (Xu et al., 2024c) mitigate color cast by adding an sRGB branch, but generating moiré-contaminated sRGB images requires an extra ISP process, which is computationally expensive (see Fig. 2(c)). These limitations highlight the need for a more efficient, single-stage raw-domain demoiréing framework that preserves both high fidelity and color accuracy while enabling fast inference. Such a framework would improve robustness, efficiency, and generalization across diverse imaging devices and real-world scenarios.
@@ 2. Related Works
Moiré patterns arise from the interference of two similar frequency patterns and frequently occur when capturing screen images, significantly degrading image quality. Demoiréing is typically formulated as an inverse problem aimed at recovering clean images from moiré-affected ones. Existing methods mainly fall into two categories:
sRGB-based demoiréing methods. Given the varying scales of moiré patterns, many methods (Yang et al., 2017a; Cheng et al., 2019; He et al., 2020; Liu et al., 2020a; Wang et al., 2021; Niu et al., 2023; Wang et al., 2023; Liu et al., 2015; Yang et al., 2017b) integrate multi-scale features to eliminate both large and small moiré patterns. Sun et al. (Sun et al., 2018) constructed a real-world dataset for LCD screen image demoiréing and introduced a multi-resolution neural network to address moiré patterns at different scales. He et al. (He et al., 2019) further enhanced multi-scale aggregation with edge-guided and attribute-aware classifiers for more robust moiré modeling. Liu et al. (Liu et al., 2018) designed a DCNN composed of coarse and fine-scale networks for refined moiré removal. Addressing limitations in training data quality, Yue et al. (Yue et al., 2020b) introduced the Moiré Removal and Brightness Improvement (MRBI) dataset and proposed an image registration algorithm based on global homography and local patch matching. Several methods also explored the frequency domain, such as wavelet-domain filtering (Liu et al., 2020b) and DCT-domain filtering (Zheng et al., 2020, 2021). Transformer-based models like Uformer (Wang et al., 2022) and Restormer (Zamir et al., 2022) have recently been applied to moiré removal. Beyond sRGB image demoiréing, Dai et al. (Dai et al., 2022) created a video demoiréing dataset and proposed a baseline model featuring implicit feature alignment and selective feature aggregation.
Raw-based demoiréing methods. Raw-domain moiré removal, unaffected by ISP demosaicing and non-linear transformations, typically yields better performance. Yue et al. (Yue et al., 2022) introduced RDNet, a raw domain demoiréing network. Further, Yue et al. (Yue et al., 2023a) extended this work by leveraging raw domain pattern distributions to develop an image-video demoiréing network, incorporating channel and spatial modulation to exploit color separation and mixing. Xu et al. (Xu et al., 2024c) combined raw and sRGB images in a demoiréing module with skip connections and an sRGB-guided ISP for color restoration. These methods highlight the superiority of raw-domain processing over sRGB methods, with less loss of high-frequency detail. However, these methods rely on a two-stage pipeline: removing moiré in the raw domain, followed by raw to sRGB mapping. While this decoupling simplifies the task, it can lead to irreversible information loss. In contrast, we propose a single-stage raw domain demoiréing network, enabling more efficient and accurate image restoration.
@@ 2.2. Learning-Based RAW Image Processing
Raw domain image restoration has gained traction due to its advantages over sRGB-based pipelines, spurring progress across several tasks (Liu et al., 2019; Wang et al., 2020; Yue et al., 2020a; Wei et al., 2020; Feng et al., 2023; Yue et al., 2023b; Xu et al., 2024a). Chen et al. (Chen et al., 2018) built the SID dataset for low-light raw images and introduced a U-Net for enhancement and denoising. Building on this, Yang et al. (Yang et al., 2023) proposed a single-stage network with feature domain adaptation to decouple denoising and color mapping for raw low-light image enhancement. Beyond low-light enhancement, Zhang et al. (Zhang et al., 2019) employed optical zoom to capture LR-HR raw image pairs for super-resolution. Xu et al. (Xu et al., 2019) designed a data generation pipeline simulating camera imaging to synthesize aligned LR-HR raw and sRGB image pairs, alongside a dual-branch CNN. Zou et al. (Zou et al., 2023) extended the raw task to HDR reconstruction and noise reduction using multi-exposure bursts. Liu et al. (Liu et al., 2023) created a smartphone-captured HDR dataset covering diverse lighting conditions. Yue et al. (Yue et al., 2023a) pioneered raw domain moiré removal, followed by sRGB conversion via a pre-trained ISP. To address multi-task challenges, Qian et al. (Qian et al., 2022) introduced a joint demosaicing, denoising, and super-resolution framework, while Sharif et al. (A Sharif et al., 2021) tackled demosaicing and denoising in sensor data. In this paper, we innovate by extracting the YCbCr from Bayer input, guiding moiré removal, and combining raw stream focuses on color mapping. This approach provides a novel direction for comprehensive raw-based image restoration.
@@ 3. Method
As shown in Fig. 3, the proposed Dual Stream Demoiréing Network (DSDNet) consists of two complementary streams, raw-to-sRGB and YCbCr-to-YCbCr, that collaboratively restore moiréd raw images. Given an Bayer input image $\mathbf{I}_{\text{bayer}}\in\mathbb{R}^{2H\times 2W}$, we first convert it into a 4-channel RGGB format, denoted as $\mathbf{I}_{\text{raw}}\in\mathbb{R}^{4\times H\times W}$, which serves as the input to the color mapping stream. Simultaneously, we average the two green channels of $\mathbf{I}_{\text{bayer}}$ and convert the result into YCbCr space using a standard transformation matrix yielding $\mathbf{I}_{\text{ycc}}\in\mathbb{R}^{3\times H\times W}$ as the input to the guidance stream. This dual-stream representation facilitates both moiré removal and color fidelity preservation. The raw and YCbCr images are independently processed by dedicated encoders, then fed into a Synergic Attention with Dynamic Modulation (SADM) module to enrich the representation with cross-domain contextual features. The integrated features are subsequently routed to two specialized streams tailored for demoiréing and color mapping. Finally, luminance-chrominance features from the YCbCr stream further guide the RGB reconstruction, improving perceptual quality.
@@ 3.1. Dual Stream Decoupling
As shown in Fig. 3, the Color Mapping Block (CMB) processes raw features for color mapping, while the Color Guidance Block (CGB) provides guidance for moiré removal and luminance correction. The detailed designs are presented in the Supplementary Material. The CMB employs a State Space Model (SSM) (Guo et al., 2024) and a proposed Gated Multilayer Perceptron (GMLP) to model the raw-to-sRGB mapping. The SSM captures row-wise correlations in a line-by-line fashion, mimicking the physical row scanning of image sensors. Prior works rely on channel attention to suppress redundancy; however, our analysis (see Fig. LABEL:fig:channel) reveals that channel redundancy is negligible and that channel attention can disrupt spatial detail preservation critical for demoiréing. To address this, we introduce GMLP, which jointly captures spatial and channel-wise patterns. Specifically, for an input feature $\mathbf{F}_{a}$, we first expand its channels, split it into two parts, apply depth-wise convolution ($\text{DConv}_{3\times 3}$) to one half, and fuse them via gated multiplication: Finally, the two parts are fused via an element-wise gating mechanism. The GMLP process is described as:
@@
[…]
Meanwhile, the CGB extracts the prior guidance features using dilated channel attention from (Xu et al., 2024c), capturing both global and local dependencies:
@@
Together, these complementary features enable moiré suppression while achieving accurate color mappings.
@@ 3.2. Synergic Attention with Dynamic Modulation
To enrich the raw-to-sRGB conversion with cross-domain contextual features, we introduce the Synergic Attention with Dynamic Modulation (SADM) module. SADM enables bidirectional interaction between raw features $\mathbf{\hat{F}}^{i}_{\text{raw}}$ and YCbCr features $\mathbf{\hat{F}}^{i}_{\text{ycc}}$ via symmetric attention mechanisms and a dynamic gating fusion pathway. Given the encoded raw features $\mathbf{\hat{F}}^{i}_{\text{raw}}\in\mathbb{R}^{c\times h\times w}$ and YCbCr features $\mathbf{\hat{F}}^{i}_{\text{ycc}}\in\mathbb{R}^{c\times h\times w}$, SADM first computes cross-domain correlations through symmetric attention pathways. In the Raw-to-YCbCr pathway, $\mathbf{\hat{F}}^{i}_{\text{raw}}$ is projected into key-value pairs ${\mathbf{K}_{\text{raw}},\mathbf{V}_{\text{raw}}}$ via a $1\times 1$ convolution ($\text{Conv}_{1\times 1}$), while the query $\mathbf{Q}_{\text{ycc}}$ is derived from $\mathbf{\hat{F}}^{i}_{\text{ycc}}$. Symmetrically, in the YCbCr-to-Raw pathway, ${\mathbf{K}_{\text{ycc}},\mathbf{V}_{\text{ycc}}}$ are obtained from $\mathbf{\hat{F}}^{i}_{\text{ycc}}$ and $\mathbf{Q}_{\text{raw}}$ from $\mathbf{\hat{F}}^{i}_{\text{raw}}$:
@@
To adaptively merge these pathways, we introduce a dynamic content-aware gating mechanism to learn spatial-wise fusion weights. The concatenated features $[\mathbf{\hat{F}}^{i}_{\text{raw}},\mathbf{\hat{F}}^{i}_{\text{ycc}}]$ are processed via two cascaded $1\times 1$ convolutions with GELU activation:
@@
where $\mathcal{G}(\cdot)$ is the dynamic gating function composed of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ ($1\times 1$ convolution), followed by GELU and Sigmoid $\sigma$ activations, producing fusion coefficients $\alpha\in[0,1]^{h\times w}$.
The final fused feature is obtained by blending the two attention outputs:
@@
This output is projected back to the original dimensionality and further refined by a feed-forward network (FFN) (Zamir et al., 2022) with a residual connection:
@@
To propagate this enriched representation, we modulate each stream using channel-spatial recalibration. Specifically, for the raw stream:
@@
(13) $\mathbf{F}^{i+1}_{\text{raw}}=\mathbf{\hat{F}}^{i}_{\text{raw}}\odot\mathbf{M}_{\text{raw}}+\mathbf{F}^{i}_{\text{fuse}},$
@@
Through this modulation mechanism, SADM enables adaptive selection and effective integration of cross-domain features, enriching the raw-to-sRGB conversion with cross-domain contextual features

### v2

@@ Abstract.
With the rapid advancement of mobile imaging, capturing screens using smartphones has become a prevalent practice in distance learning and conference recording. However, moiré artifacts, caused by frequency aliasing between display screens and camera sensors, are further amplified by the image signal processing pipeline, leading to severe visual degradation. Existing sRGB domain demoiréing methods struggle with irreversible information loss, while recent two-stage raw domain approaches suffer from information bottlenecks and inference inefficiency. To address these limitations, we propose a single-stage raw domain demoiréing framework, Dual-Stream Demoiréing Network (DSDNet), which leverages the synergy of raw and YCbCr images to remove moiré while preserving luminance and color fidelity. Specifically, to guide luminance correction and moiré removal, we design a raw-to-YCbCr mapping pipeline and introduce the Synergic Attention with Dynamic Modulation (SADM) module. This module enriches the raw-to-sRGB conversion with cross-domain contextual features. Furthermore, to better guide color fidelity, we develop a Luminance-Chrominance Adaptive Transformer (LCAT), which decouples luminance and chrominance representations. Extensive experiments demonstrate that DSDNet outperforms state-of-the-art methods in both visual quality and quantitative evaluation and achieves an inference speed 2.4x faster than the second-best method, highlighting its practical advantages. We provide an anonymous online demo at DSDNet.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754698isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computational photography
@@ 1. Introduction
Recent advancements in sRGB-domain demoiréing have been driven by deep learning methods and the availability of large, paired datasets (Cheng et al., 2019; He et al., 2020). These methods have introduced innovative strategies such as frequency-domain modeling (He et al., 2019), multiscale feature fusion (Sun et al., 2018), and moiré pattern classification (Liu et al., 2018), leading to significant improvements in moiré removal. However, these sRGB-based methods (Zheng et al., 2020; Liu et al., 2020b) suffer from two critical limitations. First, they often overlook the underlying physical processes of moiré pattern formation, especially the fact that demosaicing in ISP amplifies moiré artifacts (Yue et al., 2022), as shown in Fig. 2(a). Second, by passing through the ISP pipeline, images suffer from irreversible information loss, making restoring clean images more challenging.
In response to these limitations, recent research has shifted towards raw domain processing, which preserves the linear characteristics of the sensor signal and better aligns moiré pattern formation with the ISP pipeline. Although raw domain demoiréing frameworks have shown significant improvements over traditional sRGB methods, they still face inherent bottlenecks. Typically, these frameworks (Yue et al., 2022; Xu et al., 2024b) consist of two stages: the first stage learns the mapping from moiré to clean images, while the second stage learns the mapping from raw to sRGB, as shown in Fig. 2(b) and (c). These methods use cascaded subnetworks, where each stage enhances the output image from the previous one. With image-level data flow, only images are propagated forward, and each stage relies solely on the result of the prior stage. This design inherently increases the risk of information loss and error accumulation. Additionally, every subnetwork in each stage may incur information loss due to the downsampling operation. As a result, suboptimal performance arises from the lossy image-level data flow, where errors propagate, accumulate, and amplify across stages, ultimately failing to reconstruct texture details and color fidelity. Moreover, these two-stage pipelines (Song et al., 2023; Yue et al., 2023) face challenges in real-time inference and introduce significant computational overhead. For instance, Xu et al. (Xu et al., 2024b) mitigate color cast by adding an sRGB branch, but generating moiré-contaminated sRGB images requires an extra ISP process, which is computationally expensive (see Fig. 2(c)). These limitations highlight the need for a more efficient, single-stage raw-domain demoiréing framework that preserves both high fidelity and color accuracy while enabling fast inference. Such a framework would improve robustness, efficiency, and generalization across diverse imaging devices and real-world scenarios.
@@ 2. Related Works
Existing demoiréing methods mainly fall into two categories:
sRGB-based demoiréing methods. To handle the varying scales of moiré patterns, many methods have employed multi-scale architectures (Yang et al., 2017a; Cheng et al., 2019; He et al., 2020; Liu et al., 2020a; Wang et al., 2021; Niu et al., 2023; Wang et al., 2023; Liu et al., 2015; Yang et al., 2017b). The field has also advanced through the creation of dedicated real-world datasets (Sun et al., 2018; Yue et al., 2020) and the introduction of priors like edge guidance (He et al., 2019). Other works have explored the frequency domain using wavelet (Liu et al., 2020b) or DCT-based filtering (Zheng et al., 2020, 2021). More recently, Transformer-based models like Uformer (Wang et al., 2022) and Restormer (Zamir et al., 2022) have demonstrated strong performance. The scope has also expanded from static images to video demoiréing (Dai et al., 2022).
Raw-based demoiréing methods. Processing in the raw domain avoids artifacts from the ISP pipeline, such as demosaicing and non-linear transformations. This approach typically preserves more detail and achieves superior performance, as demonstrated in recent works for both image and video demoiréing (Yue et al., 2022, 2023; Xu et al., 2024b). However, these methods rely on a two-stage pipeline: first removing moiré in the raw domain, then mapping the result to an sRGB image. While this decoupling simplifies the task, it can lead to irreversible information loss, thereby limiting the final image quality.
@@ 2.2. Learning-Based RAW Image Processing
Raw image restoration has gained traction due to its advantages over sRGB-based pipelines, spurring progress across several tasks (Liu et al., 2019; Wang et al., 2020; Wei et al., 2020; Yang et al., 2025, 2023). Chen et al. (Chen et al., 2018) built the SID dataset for low-light raw images and introduced a UNet for denoising. Building on this, Yang et al. (Yang et al., 2025) proposed a single-stage network with feature domain adaptation for raw low-light image enhancement. Beyond low-light enhancement, Zhang et al. (Zhang et al., 2019) employed optical zoom to capture LR-HR raw image pairs for super-resolution. Zou et al. (Zou et al., 2023) extended the raw task to HDR reconstruction and noise reduction using multi-exposure bursts. Liu et al. (Liu et al., 2023) built a smartphone-captured HDR dataset covering diverse lighting conditions. To address multi-task challenges, Qian et al. (Qian et al., 2022) introduced a joint demosaicing, denoising, and super-resolution framework, while Sharif et al. (A Sharif et al., 2021) tackled demosaicing and denoising in sensor data.
@@ 3. Method
As shown in Fig. 3, the proposed DSDNet consists of two complementary streams, raw-to-sRGB and YCbCr-to-YCbCr, that collaboratively restore moiréd raw images. Given an Bayer input image $\mathbf{I}_{\text{bayer}}\in\mathbb{R}^{2H\times 2W}$, we first convert it into a 4-channel RGGB format, denoted as $\mathbf{I}_{\text{raw}}\in\mathbb{R}^{4\times H\times W}$, which serves as the input to the color mapping stream. Simultaneously, we average the two green channels of $\mathbf{I}_{\text{bayer}}$ and convert the result into YCbCr space using a standard transformation matrix yielding $\mathbf{I}_{\text{ycc}}\in\mathbb{R}^{3\times H\times W}$ as the input to the guidance stream. This dual-stream representation facilitates both moiré removal and color fidelity preservation. The raw and YCbCr images are independently processed by dedicated encoders and then fed into an SADM module to enrich the representation with cross-domain contextual features. The integrated features are subsequently routed to two specialized streams tailored for demoiréing and color mapping. Finally, luminance-chrominance features from the YCbCr stream further guide the RGB reconstruction, improving perceptual quality.
@@ 3.1. Dual Stream Decoupling
As shown in Fig. 3, the Color Mapping Block (CMB) processes raw features for color mapping, while the Color Guidance Block (CGB) provides guidance for moiré removal and luminance correction. The detailed designs are presented in the Supplementary Material. The CMB employs a State Space Model (SSM) (Guo et al., 2024) and a Gated Multilayer Perceptron (GMLP) to model the raw-to-sRGB mapping. The SSM captures row-wise correlations in a line-by-line fashion, mimicking the physical row scanning of image sensors. Prior works rely on channel attention to suppress redundancy; however, our analysis reveals that channel redundancy is negligible and that channel attention can disrupt spatial detail preservation critical for demoiréing. To address this, we introduce GMLP, which jointly captures spatial and channel-wise patterns. Specifically, for an input feature $\mathbf{F}_{a}$, we first expand its channels, split it into two parts, apply depth-wise convolution ($\text{DConv}_{3\times 3}$) to one half, and fuse them via gated multiplication: Finally, the two parts are fused via an element-wise gating mechanism. The GMLP process is described as:
@@
where $\text{SS2D}(\cdot)$ represents the 2D state space operator from (Guo et al., 2024).
Meanwhile, the CGB extracts the prior guidance features using dilated channel attention from (Xu et al., 2024b), capturing both global and local dependencies:
@@ 3.2. Synergic Attention with Dynamic Modulation
To enrich the raw-to-sRGB conversion with cross-domain contextual features, we introduce the SADM module. SADM enables bidirectional interaction between raw features $\mathbf{\hat{F}}^{i}_{\text{raw}}$ and YCbCr features $\mathbf{\hat{F}}^{i}_{\text{ycc}}$ via symmetric attention mechanisms and a dynamic gating fusion pathway. Given the encoded raw features $\mathbf{\hat{F}}^{i}_{\text{raw}}\in\mathbb{R}^{c\times h\times w}$ and YCbCr features $\mathbf{\hat{F}}^{i}_{\text{ycc}}\in\mathbb{R}^{c\times h\times w}$, SADM first computes cross-domain correlations through symmetric attention pathways. In the Raw-to-YCbCr pathway, $\mathbf{\hat{F}}^{i}_{\text{raw}}$ is projected into key-value pairs $\mathbf{K}_{\text{raw}},\mathbf{V}_{\text{raw}}$} via a $1\times 1$ convolution ($\text{Conv}_{1\times 1}$), while the query $\mathbf{Q}_{\text{ycc}}$ is derived from $\mathbf{\hat{F}}^{i}_{\text{ycc}}$. Symmetrically, in the YCbCr-to-Raw pathway, $\mathbf{K}_{\text{ycc}},\mathbf{V}_{\text{ycc}}$} are obtained from $\mathbf{\hat{F}}^{i}_{\text{ycc}}$ and $\mathbf{Q}_{\text{raw}}$ from $\mathbf{\hat{F}}^{i}_{\text{raw}}$:
@@
To adaptively merge these pathways, we introduce a dynamic content-aware gating mechanism to learn spatial-wise fusion weights. The concatenated features $\mathbf{\hat{F}}^{i}_{\text{raw}},\mathbf{\hat{F}}^{i}_{\text{ycc}}]$ are processed via two cascaded $1\times 1$ convolutions with GELU activation:
@@
where $\mathcal{G}(\cdot)$ is the dynamic gating function composed of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ ($1\times 1$ convolution), followed by GELU and Sigmoid $\sigma$ activations, producing fusion coefficients $\alpha\in[0,1]^{h\times w}$. The final fused feature is obtained by blending the two attention outputs:
@@
This output is projected to the original dimensionality and further refined by a feed-forward network (FFN) (Zamir et al., 2022):
@@
To enrich representation, we modulate each stream by channel-spatial recalibration. Specifically, for the raw stream:
@@
(13) $\mathbf{F}^{i+1}_{\text{raw}}=\mathbf{\hat{F}}^{i}_{\text{raw}}\odot\mathbf{M}_{\text{raw}}+\mathbf{F}^{i}_{\text{fuse}}.$
@@
Through this modulation mechanism, SADM enables adaptive selection and effective integration of cross-domain features, enriching the raw-to-sRGB conversion with cross-domain contextual features.